arricake:

A car goes 400 miles on 10 gallons of gas. How many miles can it go on 30 gallons?

KPalm:

What do you have to do to 10 gallons to make it 30? Do that same thing to 400 miles :)

YRJ8498:

what is 400+400+400

arricake:

@yrj8498 wrote:
what is 400+400+400
600

YRJ8498:

no try again

arricake:

@yrj8498 wrote:
no try again
1400

YRJ8498:

no try again

arricake:

@yrj8498 wrote:
no try again
i give up

YRJ8498:

400*3

KPalm:

400+400=800...plus another 400 would be?

arricake:

@kpalm wrote:
400+400=800...plus another 400 would be?
1200

YRJ8498:

Good job but do you know why we added 400 3 times?

arricake:

@yrj8498 wrote:
Good job but do you know why we added 400 3 times?
no

carmelle:

To solve this question mathematically, you have to set up a proportion. \[\frac{ 400 ~ miles }{ 10~ gallons }=\frac{ x~ miles }{ 30~gallons }\] To solve, you have to cross-multiply. \[(10)(x)=(400)(30)\] You are left with: \[10x = 12000\] Divide both sides by 12 to find your final answer. \[\frac{ 10x }{ 10 }=\frac{ 12000 }{ 10 }\] \[x=1200\] That means the car can go 1200 miles on 30 gallons of gas.

carmelle:

Divide both sides by 10* sorry typo